8 January 2015: Thursday

LDAP:

  • Three issues.  Two found and merged.  The third one is not enough to be certain and so I emailed payroll to verify the possible match.
  • VPN: Made a couple of changes per requests.

Backups:

  • Disk clones finished overnight – that makes all the clones done for this week.
  • CoM/IS groups still having issues with that disk array and VSS snapshots are failing with interesting codes.

Tags:

Leave a Reply

You must be logged in to post a comment.